OpenStudy (qaisersatti):

Vin=45V R1=1.5 R2=2 R3 6 A=2 B=0.5 please tell me Q No 6 answers time is almost over for my friend...

OpenStudy (qaisersatti):

sorry Vin is 27 V i wrote it wrong

OpenStudy (qaisersatti):

my answers are 1)Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN Vo x(12-(0.1x5x12)+5+7+6x7)/(12-0.1x5x12)=27 Vo = 2.7V 2)i1=V0/(R3-B*R1*R3) = 2.7/(12-0.1x5x12)=0.45 3)vo = vth = 2.7V 4)Rth=Vth/Isc=2.7/0.45= 6 ohm

OpenStudy (qaisersatti):

are they right